### v2

@@ Abstract
Defining pathologies automatically from medical images aids the understanding of the emergence and progression of diseases, and such an ability is crucial in clinical diagnostics. However, existing deep learning models heavily rely on expert annotations and lack generalization capabilities in open clinical environments. In this study, we present a generalizable vision-language pre-training model for Annotation-Free pathological lesions Localization (AFLoc). The core strength of AFLoc lies in its extensive multi-level semantic structure-based contrastive learning, which comprehensively aligns multi-granularity medical concepts from reports with abundant image features, to adapt to the diverse expressions of pathologies and unseen pathologies without the reliance on image annotations from experts. We demonstrate the proof of concept on CXR images, with extensive experimental validation across 4 distinct external datasets, encompassing 11 types of chest pathologies. The results demonstrate that AFLoc surpasses 6 state-of-the-art methods and even outperforms the human benchmark in locating 5 different pathologies. We further verify its generalization ability in retinal fundus image pathological lesions localization. Our approach showcases AFLoc versatilities and underscores its suitability in complex clinical environments.
@@
Pathological lesions localization plays a crucial role for advancing scientific research and disease diagnosis [1, 2, 3, 4]. In the past decade, breakthrough advancements in deep learning have significantly accelerated progress within the domain of disease localization [5, 6, 7, 8]. Nevertheless, the efficacy of these deep learning approaches is considerably constrained by their profound dependency on richly annotated training datasets [9, 10, 11, 12, 13]. Specifically, the annotation work required for localization tasks is particularly laborious, often involving precise bounding box or pixel-wise annotation of localized pathological areas. Compared to the image-level annotations (e.g., disease labels), the high-precision annotation efforts of localization tasks is more costly and thereby requires significantly more time investment.
In recent years, self-supervised pre-training [14, 15, 16] has been proposed to mitigate reliance on large annotated datasets. Initially, these methodologies acquire general visual representations via self-supervised learning from image datasets, subsequently utilizing a comparatively reduced dataset of annotated data for supervised fine-tuning of the model. This strategy enables the model to perform well on trained categories, while significantly reducing the need and cost for data labeling. Additionally, saliency methods [17, 18, 19, 20] can also reduce the annotation costs in pathological lesions localization tasks to a certain degree because they allow for the coarse localization of target categories in models trained with image-level annotations. However, in clinical practice, models may require additional training or tuning to adapt to unseen diseases [9, 21]. This challenge is particularly acute for emerging diseases [22]. In such scenarios, rapidly acquiring a dataset with expert annotations for training or adjusting the model becomes extremely difficult. Consequently, deploying these models in an open clinical environment entails substantial challenges.
@@
In this study, we present AFLoc, a vision-language pre-training model designed to alleviate the need for costly bounding boxes. AFLoc can effectively perform pathological lesions localization in chest X-ray images using brief pathologies descriptions. The overall framework of AFLoc is depicted in Fig. 1. Different from the commonly used global semantic alignment strategies [26, 25], AFLoc introduced a contrastive learning framework with a multi-level semantic alignment structure. In AFLoc, the images are processed through an encoder to generate features at three levels: shallow features, deep features, and global features. These features correspond to three different levels of embeddings extracted by the text encoder: word-level, sentence-level, and report-level. Subsequently, information from both modalities is carefully matched and aligned according to their respective granularities. The framework is capable of comprehensively aligning medical concepts from reports with image features in a shared semantic space. In the aligned semantic space, pathology can be located by matching the similarity between natural language tags/descriptions of the pathology and local images. The effectiveness of AFLoc has been validated on 4 external independent datasets, including RSNA Pneumonia [30], COVID Rural [31, 32], MS-CXR [25], and CheXlocalize [17]. The results showed that AFLoc can achieve accurate pathological lesions localization with brief descriptions for 11 types of chest pathologies and performed well in localizing unseen diseases. Moreover, AFLoc’s generalizability extends beyond chest X-rays to include pathological lesions localization in retinal fundus images, demonstrating its potential applicability in diverse clinical scenarios. This extension to retinal fundus images showcases AFLoc’s capability in handling different types of medical image, further underscoring its adaptability and efficacy in complex clinical environments.
@@ Results
We employed chest X-ray images and radiological reports to instantiate the pathological lesions localization task without the need for training on explicit labels. In our experiments, we initially engaged in pre-training utilizing the MIMIC-CXR dataset [33] and subsequently executed direct evaluations across 4 external datasets, including RSNA Pneumonia [30], COVID Rural [31, 32], MS-CXR [25], and CheXlocalize [17]. Throughout the evaluation phase, the models engendered similarity maps or heatmaps (Fig. 1) through the correlation of pathology descriptors with chest X-rays. Subsequently, thresholds were meticulously applied to each similarity map, culminating in the generation of binary segmentation maps. Furthermore, we executed comprehensive ablation studies to delineate and analyze the contributions of different components of our AFLoc.
@@
To demonstrate AFLoc versatility as a general vision-language foundation model, we investigated its adaptability in its AI-driven pathological lesions localization capability to other medical image modalities. To exemplify this ability, we chose a common ophthalmology image modality retina fundus photographs. We trained and tested the model on 26,028 retinal images. We conducted an evaluation and compared the performance with other similar models. The quantitative results are listed in Extended Data Table 3. These results demonstrated effective pathological lesions localization and advantages over other current models, achieving superior performance compared to ConVIRT and GLoRIA models (Extended Data Table 3). Examples of images are illustrated in Extended Data Fig. 6-8. These results indicated that AFLoc could adapt to a completely different medical image modality, which is valuable in clinical applications.

### v3

@@ Abstract
Defining pathologies automatically from medical images aids the understanding of the emergence and progression of diseases, and such an ability is crucial in clinical diagnostics. However, existing deep learning models heavily rely on expert annotations and lack generalization capabilities in open clinical environments. In this study, we present a generalizable vision-language pre-training model for Annotation-Free pathological lesions Localization (AFLoc). The core strength of AFLoc lies in its extensive multi-level semantic structure-based contrastive learning, which comprehensively aligns multi-granularity medical concepts from reports with abundant image features, to adapt to the diverse expressions of pathologies and unseen pathologies without the reliance on image annotations from experts. We demonstrate the proof of concept on CXR images, with extensive experimental validation across 4 distinct external datasets, encompassing 11 types of chest pathologies. The results demonstrate that AFLoc surpasses state-of-the-art methods in pathological lesions localization and disease classification, and even outperforms the human benchmark in locating 5 different pathologies. Additionally, we further verify its generalization ability by applying it to retinal fundus images. Our approach showcases AFoc versatilities and underscores its suitability for clinical diagnoses in complex clinical environments.
@@
Pathological lesions localization plays a crucial role for advancing scientific research and disease diagnosis [1, 2, 3, 4]. In the past decade, breakthrough advancements in deep learning have significantly accelerated progress within the domain of disease localization [5, 6, 7, 8]. Nevertheless, the efficacy of these deep learning approaches is considerably constrained by their profound dependency on richly annotated training datasets [9, 10, 11, 12, 13]. Specifically, the annotation work required for localization tasks is particularly laborious, often involving precise bounding box or pixel-wise annotation of localized pathological areas. Compared to the image-level annotations (e.g., disease labels), the high-precision annotation efforts of localization tasks is more costly and thereby requires significantly more time investment. The manuscript is prepared on January 20, 2024. In recent years, self-supervised pre-training [14, 15, 16] has been proposed to mitigate reliance on large annotated datasets. Initially, these methodologies acquire general visual representations via self-supervised learning from image datasets, subsequently utilizing a comparatively reduced dataset of annotated data for supervised fine-tuning of the model. This strategy enables the model to perform well on trained categories, while significantly reducing the need and cost for data labeling. Additionally, saliency methods [17, 18, 19, 20] can also reduce the annotation costs in pathological lesions localization tasks to a certain degree because they allow for the coarse localization of target categories in models trained with image-level annotations. However, in clinical practice, models may require additional training or tuning to adapt to unseen diseases [9, 21]. This challenge is particularly acute for emerging diseases [22]. In such scenarios, rapidly acquiring a dataset with expert annotations for training or adjusting the model becomes extremely difficult. Consequently, deploying these models in an open clinical environment entails substantial challenges.
@@
In this study, we present AFLoc, a vision-language pre-training model designed to alleviate the need for costly bounding boxes. AFLoc can effectively perform pathological lesions localization in chest X-ray images using brief pathologies descriptions. The overall framework of AFLoc is depicted in Fig. 1. Different from the commonly used global semantic alignment strategies [26, 25], AFLoc introduced a contrastive learning framework with a multi-level semantic alignment structure. In AFLoc, the images are processed through an encoder to generate features at three levels: shallow features, deep features, and global features. These features correspond to three different levels of embeddings extracted by the text encoder: word-level, sentence-level, and report-level. Subsequently, information from both modalities is carefully matched and aligned according to their respective granularities. The framework is capable of comprehensively aligning medical concepts from reports with image features in a shared semantic space. In the aligned semantic space, pathology can be located by matching the similarity between natural language tags/descriptions of the pathology and local images. The effectiveness of AFLoc has been validated on 4 external independent datasets, including RSNA Pneumonia [30], COVID Rural [31, 32], MS-CXR [25], and CheXlocalize [17]. The results showed that AFLoc can achieve accurate pathological lesions localization with brief descriptions for 11 types of chest pathologies and performed well in localizing unseen diseases. Moreover, this method surpasses state-of-the-art methods in disease classification capabilities. Furthermore, AFLoc’s generalizability extends beyond chest X-rays to include retinal fundus images, demonstrating its potential applicability in diverse clinical scenarios. This extension to retinal fundus images showcases AFLoc’s capability in handling different types of medical image, further underscoring its adaptability and efficacy in complex clinical environments.
@@ Results
We employed chest X-ray images and radiological reports to instantiate the tasks of pathological lesions localization and disease classification without the need for training on explicit annotations. In our experiments, we initially engaged in pre-training utilizing the MIMIC-CXR dataset [33] and subsequently conducted localization evaluations across 4 external datasets, including RSNA Pneumonia [30], COVID Rural [31, 32], MS-CXR [25], and CheXlocalize [17]. Throughout the evaluation phase, the models engendered similarity maps or heatmaps (Fig. 1) through the correlation of pathology descriptors with chest X-rays. Subsequently, thresholds were meticulously applied to each similarity map, culminating in the generation of binary segmentation maps. Furthermore, we executed comprehensive ablation studies to delineate and analyze the contributions of different components of our AFLoc. For the disease classification task, we conducted additional evaluations using the RSNA Pneumonia and SIIM-ACR Pneumothorax datasets.
@@
To demonstrate AFLoc versatility as a general vision-language foundation model, we investigated its adaptability in its AI-driven pathological lesions localization capability to other medical image modalities. To exemplify this ability, we chose a common ophthalmology image modality retina fundus photographs. We trained and tested the model on 26,028 retinal images. We conducted an evaluation and compared the performance with other similar models. The quantitative results are listed in Extended Data Table 3. These results demonstrated effective pathological lesions localization and advantages over other current models, achieving superior performance compared to ConVIRT and GLoRIA models (Extended Data Table 3). Examples of images are illustrated in Extended Data Fig. 6-8. We tested the disease diagnostic performance on chest x-ray and fundus as shown in Extended Data Table 4 and Extended Data Table 5. These results indicated that AFLoc could adapt to a completely different medical image modality, which is valuable in clinical applications.
@@ Methods
SIIM-ACR Pneumothorax Segmentation (SIIM) [37]. The SIIM dataset is designed to promote the development of models for identifying pneumothorax. It includes more than 120,000 frontal chest X-rays, each of which has been precisely annotated. Employing the data partitioning method described by [27], we conducted tests on the test set.
@@ Acknowledgments
Appendix A
Table 2: Quantatitive pathological lesions localization results (CNR) of different models on the 3 datasets, RSNA Pneumonia [30], COVID Rural [31, 32], MS-CXR [25]. Numbers within parentheses indicate 95% CI.
Methods RSNA Pneumonia Covid-Rural MS-CXR
GLoRIA 1.223 (1.179,1.271) 0.664 (0.561,0.772) 1.289 (1.149,1.430)
BioViL 0.934 (0.885,0.964) 0.776 (0.698,0.862) 1.247 (1.111,1.380)
MedKLIP 1.030 (1.013,1.050) 0.841 (0.771,0.907) 1.043 (0.962,1.131)
AFLoc 1.538 (1.496,1.578) 1.061 (0.948,1.175) 1.634 (1.494,1.775)
Table 3: Quantitative results (IoU) from the ablation study to investigate the importance of the different levels of text embeddings.
Word-level Sentence-level Report-level
RSNA Pneumonia
COVID Rural
MS-CXR
CheXlocalize
Mean
0.239 0.147 0.180 0.133 0.175
0.281 0.169 0.318 0.301 0.267
0.191 0.056 0.105 0.186 0.135
0.329 0.177 0.264 0.285 0.264
0.254 0.159 0.175 0.200 0.197
0.285 0.211 0.325 0.305 0.282
0.342 0.212 0.324 0.318 0.299
Table 4: Quantitative results of different models on the retina photograph dataset evaluated under two different evaluation metrics. Numbers within parentheses indicate 95% CI.
Methods Neovascularization Drusen Intraretinal hemorrhage Mean Metrics
ConVIRT 0.130 (0.117, 0.142) 0.120 (0.099, 0.142) 0.079 (0.074, 0.084) 0.110 (0.097, 0.123) IoU
GLoRIA 0.143 (0.131, 0.155) 0.107 (0.087, 0.131) 0.075 (0.070, 0.080) 0.108 (0.096, 0.122) IoU
AFLoc 0.321 (0.303, 0.337) 0.147 (0.123, 0.173) 0.097 (0.091, 0.103) 0.188 (0.172, 0.204) IoU
ConVIRT 0.361 (0.277, 0.439) 0.654 (0.527, 0.790) 0.255 (0.232, 0.278) 0.423 (0.345, 0.502) CNR
GLoRIA 0.498 (0.43, 0.57) 0.655 (0.516, 0.789) 0.215 (0.195, 0.237) 0.458 (0.380, 0.532) CNR
AFLoc 1.454 (1.372, 1.526) 0.929 (0.790, 1.064) 0.372 (0.349, 0.395) 0.918 (0.837, 0.995) CNR
Table 5: Comparison of Area Under the Curve (AUC) scores with other state-of-the-art methods on the zero-shot classification task, evaluating diagnostic imaging analysis performance.
Methods RSNA Pneumonia SIIM
ConVIRT 0.804 0.643
GLoRIA 0.715 0.534
BioViL 0.828 0.708
CheXzero 0.858 0.688
MedKLIP 0.869 0.892
AFLoc 0.887 0.897
Table 6: Comparative performance metrics of self-supervised and supervised learning methods on ophthalmic disease classification tasks, detailing the Area Under the Curve (AUC) for Macular Degeneration (MD), Retinopathy, Myopia, Glaucoma, Congenital Optic Disc Anomalies (CODA), and Retinal Arteriosclerosis (RAS) across varying label availability scenarios.
Methods Label fraction MD Retinopathy Myopia Glaucoma CODA RAS Mean
Self-supervised
ConVIRT 0% 0.421 0.759 0.870 0.848 0.854 0.466 0.703
GLoRIA 0% 0.914 0.592 0.774 0.926 0.803 0.636 0.774
AFLoc 0% 0.978 0.839 0.962 0.938 0.899 0.643 0.877
Supervised
AFLoc-w/o pretraining 100% 0.794 0.749 0.962 0.921 0.934 0.907 0.878
AFLoc-w/ pretraining 10% 0.955 0.916 0.998 0.996 0.999 0.999 0.977
AFLoc-w/ pretraining 100% 0.992 0.978 0.998 0.992 0.999 0.999 0.993
Figure 6: Visualizations of pathology localization by AFLoc on retinal fundus photographs of retinal hemorrhages. Black dashed polygens indicate the pathology areas marked by eye doctors, and the intensity of red in the heatmaps signifies the focus level of the model’s prediction, with deeper red indicating higher focus.
Figure 7: Visualizations of pathology localization by AFLoc on retinal fundus photographs of drusen. Black dashed polygons indicate the pathology areas marked by eye doctors, and the intensity of red in the heatmaps signifies the focus level of the model’s prediction, with deeper red indicating higher focus.
Figure 8: Visualizations of pathology localization by AFLoc on retinal fundus photographs of choroidal neovascularization. Black dashed polygons indicate the pathology areas marked by eye doctors, and the intensity of red in the heatmaps signifies the focus level of the model’s prediction, with deeper red indicating higher focus.